OpenStudy (anonymous):

What is the equation of a horizontal line passing through the point (2, 10)? y = 12 x = −8 y = 10 x = 2

OpenStudy (anonymous):

y= -1/5x+10.4. If you want it to be a bit neater the point slope form is y-10=-1/5(x-2)

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

wait how would that be one of the answers then?

OpenStudy (anonymous):

its the same answer, but in a different form

OpenStudy (anonymous):

if you were to distribute the -1/5 and then add the 10, you would end up with the same answer